Apiary Record: 25th September 2016

Putting the colonies to bed for the winter, checking that they have enough stores and that entrances are down to a minimum.   Still a bit too early for mouse guards and woodpecker protection.    Removal of Varroa treatment and varroa boards also undertaken.
Colony #6   was on a National deep, but we combined Colony #1, which was on a 14×12 several weeks ago.   We had put the 14×12 under the deep after the two colonies were united, to encourage  them to take up the stores in it,  but they thought  otherwise and pulled out the entrance block, gaining access to the main entrance and so filled it with stores rather than taking it up into the National Deep. We went through the National Deep and used any stores left to help one of the Nuc’s and then removed it, leaving just the 14×12 and a full super on top.
We also decided to under-super the colonies,   putting full supers under the brood box’s, to help them get at the stores in the cold months to come.   Removing all the Queen excluders  to clean them ready for next season.
Weather was Sunny to start with  17 Degrees C ,  but became overcast with a shower of rain towards the end.  No wind to speak of.
Hive #2   Inspected by Richard  W.     Apiguard and Eke removed, top super 3/4 full and lots of stores in 14×12 Deep. Qx also removed.
Hive #3  Inspected by Richard W.     Apiguard and Eke removed, stores a bit light but incoming Balsam nectar and ivy could fill.  check next time..  may need feeding if weather bad.
Hive #4  Inspected by Richard W.  Apiguard and Eke removed.  Full super put under brood box and another full super put above a clearer board to give (bee free) to a nuc ( 16 B).
Hive #5  Inspected by Peter Reader,  Apiguard and Eke removed,  Full super put under brood box and Qx removed.
Hive #6  Inspected by Graham Rowden,  Apiguard and Eke removed, bees did not lift stores up into National Deep, so left in 14×12, ND removed, frames of stores given to 16 C.
Hive #7  Inspected by Richard W.   Apiguard and Eke removed. Qx removed and full super left on top as this has two ND brood boxes.
Hive #7A Inspected by Peter Reader. Apiguard removed with Eke. All Shallow boxes.  Qx removed,  top super full.
Hive #8  Inspected by Graham Rowden,   Apiguard and Eke removed.  Under-supered.
Hive #15F Inspected by Graham R,  Apiguard and Eke removed.
Nuc #16B  Inspected by Graham R.  Brood chamber – National Deep. Queen still laying nice brood pattern, BAS,  Full super from hive #4 added under deep.  colony to be used again for observation hive at Blackmoor.
Nuc #16C  Inspected by Graham R.   Brood chamber -National Deep, Light on stores, so given Deep frames from Hive#6,   Rapid feeder put on with 1 ltr sugar 1:1

Apiary Records 11 September 2016

Weather fine, calm and sunny 23 degrees.  No full inspections.  First dose of Apiguard (2 weeks ago) removed and second dose applied to all except nucs.   Queen excluders removed from all hives.
Reduced to bi-weekly inspections until end of season.
Hive #2 – inspected by Richard.  Varroa > 50.
Hive #3 – inspected by Liz.  Moved final three feet into permanent site.  Varroa high > 200.
Hive #4 – inspected by Richard.  Full super removed and given to Hive #7. Varroa >100.
Hive #5 – inspected by Liz.  Varroa > 25.
Hive #6 – inspected by Liz.  Once again bees had broken out of sealed entrance in bottom 14×12 brood (plan is to force into upper brood box).  Re-sealed and taped down.  Reasonable high distribution of chalk brood on Varroa board.  Varroa  47.
Hive #7 – inspected by Richard.  Super above double brood empty. Replaced with full super from Hive # 4.  Varroa 0.
Hive #7A – inspected by Liz.  Varroa >200.
Hive #8 – inspected by Liz.  Evidence of dead wax moth.  Varroa 150.
Hive # 15F – inspected by Liz.  Several wax moth larvae removed.  Empty super removed.  Some damage to bottom super (damp?).  Bees angry – is hive location too cool and damp?  Varroa >60.
Hive #16B – on day trip to Alresford Show.
Hive #16C – queen not seen but brood seem happy.  If still not seen and no evidence of brood next time, consider combining with 16B.   Varroa 2.

Apiary Record 28 August 2016

Weather dull, windy and cool for time of year.
Hive # 2 – inspected by Peter.    Apiguard added. One empty super removed.  A lot of nectar in remaining super.  Varroa count 16.
Hive #3 – inspected by Mel.  Apiguard added.  2 supers: one empty and removed.  Varroa count 40.
Hive #4 – inspected by Mel.  Apiguard applied.  3 supers, 2 heavy.  Removed empty frames on remaining super.  Varroa count 40.
Hive #5 – inspected by Anne.  Apiguard applied.  One empty super removed.  Varroa count 3.
Hive #6 – inspected by Peter.  Unchanged since previous week as bees had broken out of sealed brood box.  Re-sealed entrancye to force into brood box above.  Apiguard applied.  Varroa count 15.
Hive #7 – inspected by Liz.  Varroa count 3. Apiguard added.
Hive #7A – inspected by Richard.  Apiguard added.  Varroa count high 70.
Hive #8 – inspected by Richard.  Apiguard added in eke above brood box. Wet box removed.  Varroa count 13.
Hive #15F – inspected by Mel.  Apiguard added. Varroa count 0.  Bees still angry.
Hive # 16 F – inspected by Richard.  Apiguard not applied as brood too small.  Varroa count 2.
Hive # 16C – inspected by Mel.  Queen found and marked White.  Apiguard not applied as brood too small.

Apiary Records 21 August 2016

Weather warm & breezy after 2 days of unseasonably cold, wet and windy conditions.  Good turnout, thanks to all!  Varroa treatment next week.  Finally…..reinstating #15H as #3.
Hive #2 – Inspected by Mel & Rowan.  Added entrance block.  No inspection of brood box or supers. Varroa count = 4
Hive #3  – Inspected by Richard.  Was #15H but almost completed its excruciatingly slow progress across the apiary (only 3 feet to go!) so now re-named.  No inspection of brood or supers.  Added entrance block.  Varroa count = 22
Hive #4 – Inspected by John.  No inspection of brood box or supers.  Entrance block added. Varroa count = 2
Hive # 5 – Inspected by Liz.  No inspection of brood or supers.  Entrance block added. Varroa count = 5
Hive # 6 – Inspected by Richard.  Need to create separate hive above 14 X 12 brood box (was Hive #1).  Swapped 14 X 12 to bottom above floor  and blocked entrance completely.  Covered with crownboard (reduced holes) and added false floor.  Replaced deep and supers.  Reduced entrance on false floor.  Varroa count = 4
Hive #7 – Inspected by Liz.  No inspection of brood or supers. Varroa count = 3
Hive #7A – Inspected by Liz.  Empty shallow at bottom moved above QE.  Frames for extraction combined into  1 super. Canadian clearer board added.  Frames to be extracted this week.  Varroa count v high at >100 (1week)
Hive #8 – Inspected by Richard.  No inspection of brood or supers.  Entrance block added. Varroa count = 18
Hive #15F – Inspected by Richard. No inspection of brood or supers. Varroa count = 3
Hive # 16 B – Inspected by Richard. Brood (all stages) on 4 frames. Queen seen. Plenty of stores. Entrance block added. Varroa count = 3
Hive # 16C – Inspected by Richard. Brood on 3 frames (2 shallows). Brood all stages. Entrance block added. Varroa = 2

Apiary Records 14 Aug 2016

Hive #1 – Combined with Hive #6
Hive #2 – Inspected by Brad.  9 frames of brood. Queen found and marked White.  2 QCS – one with dead queen and one with two dead workers.  Shortage of room to lay.  Consider adding brood frame next week?  Added super of drawn comb. Varroa =7 (1 week)
Hive #4 – inspected by Hilary.  8 frames of brood. Query sac brood last week.  Some bald brood (1 frame) found but generally a good strong colony.  Bottom & middle supers ready to extract. Top super extracted two weeks ago being filled again.  Varroa = 7 (1 week)
Hive #5 – inspected by Hilary. Brood + Half.  6 frames of brood in deep.  5 frames of brood in shallow.  Small-ish colony but healthy.  Bottom super, some capping.  Good level of stores in brood shallow.  Varroa = 0 (1 week)
Hive # 6 – inspected by Mel.  2 X 14×12.  No brood in too deep. Bottom deep not examined as bees angry.  Action for next week:  Transfer top brood box to bottom; place plastic sheet with small hole over bottom brood box (false floor) and put second brood box on top.  Varroa = 0. (1 week)
Hive #7 – inspected by Liz.  2 X deep.  Has been united with 16A.  4 frames of brood in bottom deep; 2 frames of brood in top deep. Queen found and marked white. Combined two supers to one.  Varroa  = 0 (1 week)
Hive #7A- inspected by Liz.  4 X shallows.  Brood and stores in top 3 shallows.  Bottom shallow empty.  Possibly frames of stores to extract from top 3. Varroa = >50 (1 week)
Hive #8 – inspected by Liz. Brood on 8 frames.  Super nearly empty. Stores in Brood Box. Varroa = 50 (1 week)
Hive #15F – inspected by Richard.  Brood on 6 frames (bas). Bees angry.  Top super half full.  Most stores in super above brood box.  Varroa = 2(1 week)
Hive #15H – inspected by PG.  Deep X 2. Queen seen and re-marked mblue.  Deep1 not inspected.  Brood on 6 frames in Deep2.   Super empty. Moved another 3 feet towards Hyive 3 position. Varroa = 15 (1 week)
Hive 16B – inspected by Richard/PG.  Nuc transferred to national deep on 24/7. Brood on 5 frames.  Queen found and marked white.  Drone trap left in as full of stores.  Varroa = 2
Hive 16C – inspected by Richard/PG.  Nuc created 26 June.  Several shallow frames in use so left in.  4 deep frames of foundation swapped for empty shallows. Varroa = 1 (1 week)

Apiary Records – Sunday 24 July 2016

Weather breezy, warm & cloudy.  Starting to rain.  Identified supers ready for extraction next week (porter bee escapes to be added on Thursday).
Hive#1 – Not checked, poss virgin queen
Hive#2 – Not checked, poss virgin queen
Hive#4 – Inspected by Brad.  Brood on 9 frames, question whether sufficient room. 4 supers, 2 ready for extraction.  Swap brood frame for foundation at next inspection.  Queen could do with re-marking. Add super next week.  Some sac brood.
Hive#5 – Inspected by Mel.  Brood + half.  Brood on 1 frame in deep.  Stores and lots of pollen.  Supers not ready for extraction. Super added last week not started to be filled.  Feisty bees.
Hive#6 – not checked.
Hive#7  – inspected by Liz.   Searched for new queen but not seen.  Temper good so poss queenright.  Believe queen in bottom but QE under top deep.  Bottom super ready to extract next week.
Hive #7A – not checked.
Hive #8 – inspected by Richard.  Queen not seen. Brood on 7 frames.  Super not ready for extraction.  Plenty of room.
Hive#15F – inspected by Peter.  QC knocked down.  Queen seen.  Angry bees.  Second super added. Brood on 7 frames.
Hive#15H – inspected by Richard & Mel.  2 X deep (no supers last inspection).  Added super.  Top super ready to extract next week.  Moved a further 3 feet towards Hive 3 site.
Hive 16 B – inspected by Brad.  Very strong Nuc. Brood on 5 frames – no room.  Needs hiving.
Hive 16C – inspected by Richard.  Nuc.  Brood on 3 framnes. Needs super frames replacing with deeps after brood hatched.

Apiary Records: Tuesday 19th July

Was meant to be SBI  recheck on colonies to give clearance from EFB.   Due to SBI being off sick it was cancelled at the last minute,  but as we had delayed going through the colonies on Sunday we had to check them over.
Weather :-  HOT  and Sunny,  Little wind,  24 degrees.
Hive#1    not gone through as poss Virgin Queen
Hive#2    not gone through as poss virgin Queen
Hive#4    Inspected by Richard and Liz.   Brood Chamber:   National Deep.  QNS, no Qc’s. Brood and new eggs, room/stores ok.   4 supers,  3 ready to take off.  –  lots of play cups but none charged.  Temper 7
Hive#5   Inspected By Richard and Liz.    Brood chamber: Deep and Half.   QNS,   3 hatched Qc’s.  1 deep frame of brood/eggs . lot of drones – first time gone through for a month.  Temper 7
Hive#6  Not gone through as poss virgin Queen.
Hive#7  Not gone through
Hive# 7A    Inspected by  Richard and Liz,   Brood chamber:  4 Shallows ( Peter Street’s Method)                QNS,  Several play cups, not charged,  5 f bas in top super, room and stores ok. temper 8
Hive#8     Inspected by Richard and Liz.    Brood chamber: Deep.   QNS,  no Qc’s.  2 f BAS,  stores ok but room very limited,  Q laying right up to the frames .   2 supers, both near full, needs another next visit.  Temper  9.
Hive# 15 F,   Inspected by Richard and Liz,  Brood chamber: Deep.    QS,   1 QC,  Knocked down. 6 f BAS,  stores /room ok.  1 super( nearly full) –  Needs another Super next visit. Temper 9
Hive#15 H,    Inspected by Richard and Liz,  Brood chamber: Deep.   QNS,  3 Charged Qc’s knocked down. 6 F BAS with new eggs.  lack of room to lay so made up another Deep and moved another 3 ft towards Hive 3 slot.   Temper: 9.
Nuc’s  16B and 16C  not gone through.

Apiary Record : 10th July 2016.

Still under EFB shutdown until being rechecked by the SBI on the 19th July.  So quick check through, making sure no colonies were about to swarm and that they had enough room,stores and that they had no sign of any disease.
Weather:   Sunny after earlier rain/drizzle.  19 degrees C,  Breezy.
Hive#1   Inspected by Liz,    Brood Chamber  14 x 12,  QNS,  No Qc’s,  No Brood, No eggs, stores/room ok.   1 super  75% full,   Queenless..  Introduced a frame with 2 QC’s from 15H( soon to be the new Colony 3).  Drone Trap removed,  far too many drones in the hive. leave 4 weeks.  Temper : 9
Hive#2   Not Checked as sealed QC left on f2 on 3/7/16    will be checked on 19/7 by SBI
Hive#4  Inspected by Liz.   Brood chamber -National Deep.  QNS,  2 Qc’s Brood  AS,  and new day old eggs,  2 Supers  nearly full, stores/room ok.   –  Knocked down both charged Qc’s.  Temper: 7
Hive#5  Not gone through as poss virgin Queen.
Hive#6  Inspected by Liz,    Brood chamber,  Deep and a Shallow.   QNS, possibly swarmed.        4 charged Qc ‘s   Brood  5D, 4S,  2 supers, top full, –  knocked down 2 Qc’s leaving 1 sealed, and 1 charged but open, on nframe with red pin.  No eggs seen at all.-  leave for 4 weeks.                   Temper:  6.5
Hive#7  Inspected by Liz,   Brood chamber: Deep.   QNS,  no Qc’s, stores/room ok,  2 supers.     new eggs in drone comb only, so may be a drone laying queen, Combined with Nuc 16A as 2 Sealed Qc’s on f5.
Hive#7A   Inspected by Richard,    Brood chamber:  4 Shallows.   QNS, no Qc’s.  7 frames of brood, all in top super. some 6 -7 day old larva,  no new eggs.  Q gone?  Dirty comb in bottom supers needs replacing.
Hive#8    Inspected by Richard,    Brood chamber : Deep.   QNS, 2 charged Qc’s,  6 f BAS, new eggs, 2 supers, top full.  stores/room ok,   Charged Qc’s taken down.  lots of bees looks like they may swarm.
Hive#15F  Inspected by Liz.     Brood chamber: Deep.  QS,  no Qc’s.  6 f BAS,  stores/room ok,     1 super,  –  Knocked down several play cups.
Hive#15H   Inspected by Richard.   Brood chamber: Deep.   QS,  2 QC’s ,  7 f BAS and new eggs.  Stores/room ok,  1 super . Part filled,-   Frame with 2 Qc’s given to Hive 1.    Temper:  9
Nuc # 16A   Combined with Hive#7.   but no Q found, 2 sealed QC’s  left.
Nuc# 16B   Not gone through – though bees bringing in pollen
Nuc# 16C  Not gone through –  bees flying

Apiary Record: 3 July 2016

Apiary continues to be in EFB lockdown.  6-week inspection due 10 July.
Hive #1 – not checked
Hive #2 – continues to set down charged queen cells.  9 frames of sealed brood, some older larvae but no eggs seen.  Removed 10 charged QCs, left 2.   Drawing top super, some capped stores.
Hive #4 – 9 frames of brood.  15+ play cups, 3 with eggs but not charged.  Big colony.  Some bald brood and sac brood.
Hive #5 – not checked.
Hive #6 – Took down 2 queen cells leaving 1 sealed 1 charged.  Marked frame with red pin.  No eggs.   Brood on 5 deep frames and 4 shallows.
Hive #7 – 1 day old eggs in drone comb only.
Hive #7A –  Sealed brood but no eggs/larvae in any shallows below QX.      Large number of dead drones on top of QX.  Eggs/larvae/sealed brood in super.  Removed QX.  Bees grumpy.
Hive #8 – new eggs 1-3 days.  Calmed down and temper fine. Super half full.
Hive #15H – continues its weekly migration towards Hive #3 site.  Queen seen.  5 queen cells – all removed.  Queen seen.  Better temper.
Hive #15F – 6 frames of brood, 3 stores.  Whole frame of new eggs.  Good temper.
Hive #16A – Nuc created from Hive #4.  Queen found and marked white (NB one worker got in the way and also marked white!).  Transferred all frames to deep, plus frame of stores.  Some sealed cells opened (random pattern) and varroa seen.
Hive #16B – Nuc created from Hive #4.  Brood in all stages, healthy development.  Queen seen but disappeared before marker pen located.

Apiary Record: 26th June. From memory….

To follow the EFB  “lockdown” to the letter, we had to leave all the record sheets in the container so a rough guide from memory (known to be unreliable ).
Hive #1   Not checked as poss Qc / Virgin Queen.
Hive #2  Checked and QS,   BAS with new eggs.   lots of play cups taken down but none charged. top super full ready to take off when we can.
Hive #3   Culled
Hive#4   Checked and all  OK.   Top 2 supers full,  3rd filling.
Hive#5   Not checked  as  poss Virgin Queen.
Hive #6  Checked and QS ,  BAS and eggs,   6 Qc’s  some sealed,  another Nuc made up from a frame with two nice Qc’s on , along with a frame of brood and a frame of stores.  The remaining Qc’s were taken down.  Very nice tempered colony.
Hive #7  Not Checked as poss VQ.
Hive #7A  Not checked as poss VQ.
Hive #8  Not checked as poss VQ.
Hive #9  Warre Hive   – Culled as a precaution, due to being a very small colony and thought to be more susceptible to EFB and not making up.
Hive#10  Top Bar Hive –  Also a very small colony and culled as per Warre hive.
Hive #15F  Ex  Nuc, now in Deep, with super added.  building up nicely.
Hive #15H  Ex Nuc,  also building up, QS. a few playcups, 5f BAS , 1 super – moved another meter towards No3 space.
Nuc #16A   From Hive 4,  QNS,  2 play cups, waxmoth larva killed. making up nicely, 4f BAS lots of eggs.  Temper 10.   2 varroa in 4 weeks.
Nuc #16B  From Hive 4, 19/6  left as poss VQ
Nuc #16C  From Hive 6  26/6  put into wooden Nuc, furthest from Apiary on stand in ferns,